OpenStudy (anonymous):

WILL GIVE MEDALS Graph the solution to the equation 7/x -10/x^2=1

OpenStudy (anonymous):

@Hero can you help

OpenStudy (anonymous):

\[\frac{ 7 }{ x }-\frac{ 10 }{ x ^{2} }=1\]

OpenStudy (campbell_st):

multiply every term by x^2 and you get 7x - 10 = x^2 or x^2 -7x + 10 = 0 hope you can solve the quadratic by factoring

OpenStudy (anonymous):

can you help me solve it so i can put it in a calculator

OpenStudy (campbell_st):

ok... so what are the 2 factors of 10 that add to -7... both factors are negative...

OpenStudy (anonymous):

-2 and -5

OpenStudy (campbell_st):

thats it so the factored form of the equation is (x -2)(x - 5) = 0 so you need to solve x - 2 = 0 and x - 5 = 0 to find the solutions... graphically if you graph y = x^2 - 7x + 10 the solutions are where the curve cuts the x-axis here is an easy to use website that allows you to graph the curve https://www.desmos.com/calculator

OpenStudy (anonymous):

thats the site i use its great thanks so much